OpenStudy (anonymous):

PLEASE HELP and explain FACTOR y^4-81

OpenStudy (raden):

so, y^4 - 81 = (....)^2 - (....)^2

OpenStudy (anonymous):

y and 9?

OpenStudy (raden):

y^2 not be y^4, but right 9^2=81

OpenStudy (anonymous):

so what do you do next

OpenStudy (raden):

it should be y^4 - 81 = (y^2)^2 - (9)^2 next use the property : a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

ok i think i got it, how about 6a^4+13a^2-5

OpenStudy (raden):

show ur answer

OpenStudy (anonymous):

(y^2+9)(y-3)(y+3)

OpenStudy (raden):

ok, correct...

OpenStudy (raden):

for the 2nd question : 6a^4+13a^2-5 = (2a^2+5)(3a^2-1) bonus for u :)